# Transformer Block Coupling and its Correlation with Generalization in LLMs

### v2

@@ A.5 Experimental and Implementation Details
A.6 Additional Experimental Results
@@
Figure 22: Coefficient of variation of layer-wise expodistance. Variation of layer-wise expodistance (Section 3.3) computed over 1,200 prompts from the HuggingFace Open LLM Leaderboard datasets (Section 4.2) on a suite of untrained LLMs (Appendix A.1).
Figure 23: Various Measurements of Representations. All measurements were made on 100 prompts taken from the WikiText 2 datsets. (Row 1) Norm of hidden representations as a function of layer depth. (Row 2) Line Shape Score (LSS) of the hidden trajectories as a function of layer depth. (Row 3) Mean equidistance of contiguous hidden trajectories as a function of depth. (Row 4) Entropy of logit vectors as a function of depth. Noted in most plots is a line where the behaviour of the measurement drastically changes.

### v3

@@ A.6 LLM Evaluation and Implementation Details
The computational complexity for the metrics utilized in this paper are as follows:
Coupling. Coupling requires computing the Jacobians for each transformer block, and so a forward pass and backward pass required (note that we compute the Jacobians on a block level). Once the Jacobians are obtained, it requires computing a truncated singular value decomposition of each Jacobian. The time complexity of computing the truncated SVD of rank $k$ for a $d\times d$ matrix is $\mathcal{O}(d^{2}k)$, where $k<<d$. Computing $A$ from the SVDs then has time complexity $\mathcal{O}(k^{3})$, so the asymptotic time complexity of computing the coupling score between two connections is $\mathcal{O}(d^{2}k)$.
LSS. For each trajectory, the time complexity of computing the LSS is $\mathcal{O}(Ld)$ where $L$ is the number of layers and $d$ is the hidden dimension. Therefore, for a prompt containing $T$ tokens, the total time complexity for each prompt is $\mathcal{O}(TLd)$ (in addition to a single forward pass of the model).
Expodistance. Similarly, computing the expodistance of a single trajectory has time complexity $\mathcal{O}(Ld)$. Therefore, for a prompt containing $T$ tokens, the total time complexity for each prompt is $\mathcal{O}(TLd)$ (in addition to a single forward pass of the model).
A.7 ViT Training Details
For further investigation of coupling in transformers, we train 64 Vision Transformers (ViTs) following the default configurations of DEiT training (Touvron et al., 2021) on CIFAR10 (Krizhevsky, 2009). For a fixed ViT architecture with embedding dimension 192, depth of 12 layers, and 3 attention heads, we vary the weight decay $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$ and stochastic depth rate $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$. Optimization uses ADAM optimizer with $45e-$ 4 learning rate and cosine scheduler for 500 epochs with 30 epochs of linear warmup. Training proceeds with data mixup using $\alpha=0.8$ Optimization proceeds on 4 NVIDIA Tesla T4 GPUs with 128 batch size and data parallelization, with total training time being approximately 2 hours.
A.8 Additional Experimental Results
Figure 8: Exponential fit between coupling and performance.
Figure 9: Plotting (1) the number of tokens used to train each model against its mean depth-wise coupling score against and (2) the number of training tokens against its LLM Huggingface Benchmark score.
Figure 10: Utilizing the Easy2Hard Benchmark dataset Ding et al. (2024), the difficulty score of a prompt is plotted against the mean depth-wise coupling over that prompt. Additionally, plotted on the right is the prompt length (in tokens) plotted against the mean depth-wise coupling on that prompt.
Table 3: $R^{2}$ for Various Models (across different datasets)
Model $\bm{R^{2}}$
Llama2-7B 0.5484
MPT-30B 0.9629
Gemma-7B 0.7832
Mistral-v0.1 7B 0.5876
Subset $R^{2}$ $p$-value
All models $0.75$ $1.56\times 10^{-6}$
$7$B models only $0.55$ $0.023$
Score $>45$ $0.39$ $0.023$
Table 4: Summary of $R^{2}$ and $p$-values for different subsets of models for the correlation between coupling and performance.
Figure 11: Coupling against Stochastic Depth Rate. Plots are generated for each weight decay in $\{0.005,0.01,0.02,0.04,0.06,0.08,0.1,0.2\}$.
Figure 12: Transformer Block Coupling for Fixed Weight Decay. Coupling plotted against accuracy for 64 ViTs with stochastic depths $\{0,0.025,0.05,0.075,0.1,0.15,0.2,0.3\}$.
Figure 16: Evolution of Hidden Trajectories Throughout Training. Principle component visualizations of the hidden trajectories in Pythia 12B at training checkpoints 1, 4000 and 143000 on the prompt: What is the capital of France? The capital is.
@@ A.8 Additional Experimental Results
Figure 24: Various Measurements of Representations. All measurements were made on 100 prompts taken from the WikiText 2 datasets. (Row 1) Norm of hidden representations as a function of layer depth. (Row 2) Line Shape Score (LSS) of the hidden trajectories as a function of layer depth. (Row 3) Mean equidistance of contiguous hidden trajectories as a function of depth. (Row 4) Entropy of logit vectors as a function of depth. Noted in most plots is a line where the behaviour of the measurement drastically changes.
@@
Figure 34: Perturbation Experiments with Llama-3 8B. The last token embedding is perturbed with various noise levels, and compared with the true embedding at the first and last layers. The trend shows that at small noise levels, cosine similarity with the true embedding remains somewhat high, and is significantly lower at the first layer.